OpenStudy (anonymous):

y varies inversely as twice x. When x = 3, y = 6. What is the value of y when x = 9?

Parth (parthkohli):

\( \color{Black}{\Rightarrow y = k \Large \times {1 \over x}}\)

Parth (parthkohli):

Find the value of \(k\) first.

OpenStudy (anonymous):

what is k? how?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 6 = k\times{ 1\over 3} }\)

OpenStudy (anonymous):

is the answer 2?

Parth (parthkohli):

No no

Parth (parthkohli):

First find the value of k

OpenStudy (anonymous):

yeah-- k=18--- than y =2

Parth (parthkohli):

\( \color{Black}{\Rightarrow y = 18 \times 9 }\) We already know what x is.

Parth (parthkohli):

Oops

Parth (parthkohli):

Yes, it's 2.

OpenStudy (anonymous):

XD thankyou